how to find the derivative of 2x-9/(3x+4)^2


1
Expert's answer
2021-01-07T13:44:17-0500

"(2x-\frac{9}{(3x+4)^2})' = (2x-9*(3x+4)^{-2})' = 2+18*(3x+4)^{-3}*(3x)' =\\= 2+\frac{54}{(3x+4)^3}"